One number is five more than another and their sum is three less than three times the smaller. Find the numbers.
Answer by nycsharkman(136)

One number is five more than another and their sum is three less than three times the smaller. Find the numbers.

Let x = smaller number
Let x + 5 = bigger number
smaller number PLUS bigger number = 3 times smaller number minus three.
Here's what the statement looks like as an equation:
x + x + 5 = 3x - 3
Now, solve for x.
2x + 5 = 3x - 3
2x - 3x = -3 -5
-x = - 8
x = 8...smaller number
x + 5 becomes 8 + 5 = 13...bigger number
The smaller number is 8 and and bigger number is 13.